The rate constant of a certain first order reaction is doubled when the temperature is increased from
27 °C to 37 °C. Calculate the energy of activation and Arrhenius factor for the reaction.
1
Expert's answer
2020-04-10T13:36:37-0400

As "K=Ae^{-\frac{E_a}{RT}}"

So, at "T=27\u00b0C=300K"

"K=Ae^{-\frac{E_a}{300R}}"

At "T=37\u00b0C=310K"

"2K=Ae^{-\frac{E_a}{310R}}"

Dividing the two equations, we get

"2=e^{\frac{E_a}{300R}-\frac{E_a}{310R}}"

"log2=\frac{E_a}{300R}-\frac{E_a}{310R}"

"E_a=53582.8986J"

Now, "K=Ae^{-\frac{E_a}{300R}}=4.6779\times10^{-10}A"

Hence, Arrhenius factor "A=0.2137\times10^{10}K"